OpenStudy (anonymous):

a particle travels along the x axis with its position as a function of time of x(t) = 7t^4 + 24t^3 +3t where all quantities are SI units. calculate the velocity and acceleration as a function of time.

OpenStudy (anonymous):

velocity=dx/dt \[\frac{ dx }{dt }=28t ^{3}+72t ^{2}+3\] and\[ acceleration=\frac{ d ^{2}x }{dt ^{2} }\] \[\frac{ d ^{2}x }{dt ^{2} }=84 t ^{2}+144 t\]

OpenStudy (anonymous):

thank you very much!

OpenStudy (anonymous):

yw